OpenStudy (anonymous):

The radius of a cylinder is decreasing at the rate of 4cm/s, while the height is increasing at the rate of 2cm/s. Find the rate of change of the volume when the radius is 2cm and the height is 6cm. , while the height is increasing at the rate of 2cm/s. Find the rate of change of the volume when the radius is 2cm and the height is 6cm.

OpenStudy (anonymous):

the volume decreases by a rate of 276.32cm^3/s ...note value of pi approx to 3.14

OpenStudy (anonymous):

dv/dt = d(pi r^2h)/dt

OpenStudy (anonymous):

This question was posted before by same person, was something wrong with the answer?

OpenStudy (anonymous):

isnt my answer correct...???

OpenStudy (dumbcow):

\[\frac{dV}{dt} = \frac{dr}{dt}*\frac{dV}{dr} + \frac{dh}{dt}*\frac{dV}{dh}\]